# Kinetic Theory of Quasiparticles, Retarded Correlators and Hydrodynamics

### v1

@@ VI Conclusion
Kinetic theory provides a robust framework for investigating macroscopic and microscopic properties of many-body systems. This capability stems from the Boltzmann equation, which treats particles as probabilistic entities, allowing for the incorporation of every relevant detail of the system. The central probabilistic variable is the one-particle distribution function, defined as the number of particles within a given space-time and momentum interval. The dynamics of this function are governed by the Boltzmann equation. Upon solving the Boltzmann equation, the macroscopic properties of the system can be derived from specific moments of the distribution function. Notably, kinetic theory serves as an alternative approach to solving the intricate Navier-Stokes equations.
@@
With this advantage in mind, we examine the collective behavior of a system with fixed mass values. To the best of our knowledge, this work represents the first attempt to extend prior studies to the case of a massive relativistic plasma. Within the framework of Relaxation Time Approximation (RTA) kinetic theory and assuming an equilibrium state described by the Maxwell-Boltzmann distribution, we solve the Boltzmann equation for small fluctuating perturbations. This approach permits us to derive the two-point retarded correlation functions using variational formalism. These retarded two-point functions are crucial for studying bulk properties, as they contain essential physical insights. For instance, functions with poles yield valuable information about propagating modes, while transport coefficients can be explicitly determined by comparing the low-frequency limit of these functions with their counterparts in general hydrodynamic relations.
Our analysis reveals that the constituent mass shifts the hydrodynamic diffusive pole toward the unstable region, i.e., $\text{Im}\,\omega\geq 0$. This occurs due to the reduction of the mass-induced diffusion coefficient as the mass increases. This behavior holds for small mass-to-temperature ratios ($x=m/T$), while for significantly large $x$, no meaningful diffusive behavior is observed. Notably, at $x\simeq 1.2342$ and beyond, the diffusive poles exhibit instability. In the energy-momentum sector of perturbations, we have computed the complete spectrum of retarded Green’s functions and verified the fulfillment of the Ward identity. The relevant transport coefficients have been derived up to third order in momenta and frequencies. A key observation in the sound channel is that all hydrodynamic modes are purely imaginary and susceptible to instability. In contrast, the structure of hydrodynamic modes in the shear channel (comprising non-hydrodynamic and diffusive modes) remains unchanged, though the mass modifies the coefficients in the dispersion relation. Importantly, the mass preserves the branch-cut structure, and for $k\tau_{eq}\leq\pi/2$, the hydrodynamic modes lie above the branch cut in the strong coupling regime, indicating the emergence of collective modes. Similar to the diffusive charge sector, the mass displaces the lowest-lying modes toward instability. Likewise, the second law of thermodynamics and stability of modes in the shear channel put constraints on the accessible $x$ regions. These results have shown that stable thermodynamic states for massive particles are attainable for small mass-to-temperature ratios.
@@
This study underscores the critical role of constituent mass in relativistic plasmas, particularly in destabilizing hydrodynamic modes. Several promising extensions of this work can be pursued. For instance, investigating a weakly coupled quark-gluon plasma with realistic thermodynamic densities and an appropriate equation of state could provide deeper insights into the role of mass in more physical scenarios. Moreover, exploring plasmas with spatially or energy and momentum-dependent mass profiles may reveal further dynamical effects, particularly near phase transitions. We follow the strategy of “$x$” expansion, but one can perform the integrals numerically and examine the collective behavior at each mass-to-temperature value. Given the significance of mass effects near critical points, this framework could be instrumental in studying the QCD phase diagram, where hydrodynamic instabilities may play a crucial role.
@@ Appendix A Useful Integrals
By assuming the following relations [58]
@@
$K_{n+1}(x)=K_{n-1}(x)+\frac{2n}{x}K_{n}(x),\qquad\frac{\partial K_{n}(x)}{\partial x}=-\frac{K_{n-1}(x)+K_{n+1}(x)}{2},$ (72)
we obtain the following results
@@
$\int\,\frac{dp}{2\pi^{2}T_{0}}\,f_{eq}(p)=\frac{1}{2\pi^{2}}xK_{1}(x),$
$\int\,\frac{dp}{2\pi^{2}T_{0}}\,p_{0}^{2}\,f_{eq}(p)=\frac{T_{0}^{2}}{2\pi^{2}}x^{2}\left(xK_{1}(x)+K_{2}(x)\right),$
@@ Appendix B Definition of variables given in the energy-momentum sector
$b\equiv\frac{1}{ik\tau_{eq}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,\vec{k}),$
$c\equiv 1-\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,\vec{k}),$
$d\equiv 1-\frac{1}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,\vec{k}).$ (74)
@@
$\hskip 28.45274pt+2\Gamma^{0}_{03}(\omega,\vec{k})\frac{1}{ik\tau_{eq}}\mathcal{I}_{1}(v,\omega,\vec{k})+\Gamma^{0}_{33}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,\vec{k})+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\frac{\mathcal{I}_{4}(v,\omega,\vec{k})}{i\tau_{eq}\omega}\bigg{\}},$
$\mathcal{E}_{3}\equiv-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{1}{ik\tau_{eq}}\mathcal{I}_{1}(v,\omega,\vec{k})+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{5}(v,\omega,\vec{k})$
$\hskip 28.45274pt+2\Gamma^{0}_{03}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,\vec{k})+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{6}(v,\omega,\vec{k})\bigg{\}},$
$\mathcal{I}_{1}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\vec{v}(\vec{p})\cdot\vec{k}}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=1+\frac{i\,\omega\,\tau_{eq}-1}{2i\,\tau_{eq}\,k\,v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)},$
$\mathcal{I}_{2}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=-\frac{i}{k\,\tau_{eq}}\mathcal{I}_{1}(v,\omega,\vec{k}),$
$\mathcal{I}_{3}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\,\omega\,v_{z}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\frac{i\,\omega(1-i\tau_{eq}\,\omega)}{k^{2}\,\tau_{eq}}\,\mathcal{I}_{1}(v,\omega,\vec{k}),$
$\mathcal{I}_{4}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\,\omega\,v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=-\frac{i\tau_{eq}\omega}{2}\left(\frac{1-i\tau_{eq}\omega}{(\tau_{eq}\,k)^{2}}+\frac{(1-i\tau_{eq}\omega)^{2}+\tau_{eq}^{2}k^{2}v^{2}}{2v(i\tau_{eq}\,k)^{3}}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\right).$
$\mathcal{I}_{5}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{3}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\frac{(1-i\tau_{eq}\omega)^{2}}{(ik\tau_{eq})^{3}}+\frac{v^{2}}{3ik\tau_{eq}}-\frac{(1-i\tau_{eq}\omega)^{3}}{2vk^{4}\tau_{eq}^{4}}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)},$
$\mathcal{I}_{6}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{7}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{4}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{8}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{2}(\vec{p})v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\int\frac{d\Omega}{4\pi}\frac{v_{z}^{2}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{9}(v,\omega,\vec{k})\equiv\int\frac{d\Omega}{4\pi}\frac{v_{x}^{2}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$

### v2

@@ VI Conclusion
Kinetic theory provides a robust framework for macroscopic and microscopic properties of many-body systems. This capability stems from the Boltzmann equation, which treats particles as probabilistic entities, allowing for the incorporation of every relevant detail of the system. The central probabilistic variable is the one-particle distribution function, defined as the number of particles within a given space-time and momentum interval. The Boltzmann equation governs the dynamics of this function. Upon solving the Boltzmann equation, the macroscopic properties of the system can be derived from specific moments of the distribution function. Notably, kinetic theory serves as an alternative approach to solving the intricate Navier-Stokes equations.
@@
With this advantage in mind, we examine the collective behavior of a system with fixed mass values. To the best of our knowledge, this work represents the first attempt to extend prior studies to the case of a massive relativistic plasma. Within the framework of Relaxation Time Approximation (RTA) kinetic theory and assuming an equilibrium state described by the Maxwell-Boltzmann distribution, we solve the Boltzmann equation for small fluctuating perturbations. This approach permits us to derive the two-point retarded correlation functions using the variational formalism. These retarded two-point functions are crucial for studying bulk properties, as they contain essential physical insights. For instance, functions with poles yield valuable information about propagating modes, while transport coefficients can be explicitly determined by comparing the low-frequency limit of these functions with their counterparts in general hydrodynamic relations. In the current-current correlation functions, we obtain the collective modes as well as the diffusion constant for both small and large $x$ values. In the energy-momentum sector of perturbations, we have computed the complete spectrum of retarded Green’s functions and verified the fulfillment of the Ward identity. The relevant transport coefficients have been derived up to third order in $(\omega,k)$ for extreme $x$ limits. A key observation in the sound channel is that all hydrodynamic modes are purely imaginary. In contrast, the structure of hydrodynamic modes in the shear channel (comprising non-hydrodynamic and diffusive modes) remains unchanged, though the mass modifies the coefficients in the dispersion relation. Importantly, the mass preserves the branch-cut structure, and for $k\tau_{eq}\leq\pi/2$, the hydrodynamic modes lie above the branch cut in the strong coupling regime, indicating the emergence of collective modes.
@@
This study underscores the critical role of mass in relativistic plasmas. Several promising extensions of this work can be pursued. For instance, investigating temperature or energy-dependent mass profiles with an appropriate equation of state could provide deeper insights into the role of mass in more physical scenarios. This is valuable because near the phase transition points, mass serves as an order parameter running with temperature or any other parameters. Given the significance of mass effects near critical points, this framework could be instrumental in studying the QCD phase diagram, where hydrodynamic instabilities may play a crucial role. We follow the strategy of “$x$” expansion, but one can perform the integrals numerically and examine the collective behavior at each mass-to-temperature value.
@@ Appendix A Useful Integrals
In this part, we are going to provide helpful integral relations that were used in previous sections. The following relations can be derived easily [58]
@@
$K_{n+1}(x)=K_{n-1}(x)+\frac{2n}{x}K_{n}(x),\qquad\frac{\partial K_{n}(x)}{\partial x}=-\frac{K_{n-1}(x)+K_{n+1}(x)}{2}.$ (82)
Afterward, we obtain the following results
$\int\,\frac{dp}{2\pi^{2}T_{0}}p^{6}\,f_{eq}(p)=\frac{15T_{0}^{6}}{2\pi^{2}}x^{4}K_{4}(x),$
@@
$\int\,\frac{dp}{2\pi^{2}T_{0}}p^{6}\,p_{0}^{2}\,f_{eq}(p)=\frac{15T_{0}^{8}}{2\pi^{2}}x^{4}\left(7xK_{3}(x)+(56+x^{2})K_{4}(x)\right),$
$\int\,\frac{dp}{2\pi^{2}T_{0}}p^{4}\,p_{0}^{2}\,f_{eq}(p)=\frac{3T_{0}^{6}}{2\pi^{2}}x^{3}\left(5xK_{2}(x)+(30+x^{2})K_{3}(x)\right),$
@@
$\int\,\frac{dp}{2\pi^{2}T_{0}}\,f_{eq}(p)=\frac{1}{2\pi^{2}}xK_{1}(x),$
$\int\,\frac{dp}{2\pi^{2}T_{0}}\,p_{0}^{2}\,f_{eq}(p)=\frac{T_{0}^{2}}{2\pi^{2}}x^{2}\left(xK_{1}(x)+K_{2}(x)\right),$ (83)
where $x=m/T_{0}$ and $p_{0}=\sqrt{p^{2}+m^{2}}$. Likewise, we benefit from the following integrals
@@ Appendix B Definition of variables given in the energy-momentum sector
$b\equiv\frac{1}{ik\tau_{eq}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,k),$
$c\equiv 1-\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{1}(v,\omega,k),$
$d\equiv 1-\frac{1}{i\tau_{eq}\omega(\varepsilon_{eq}+P_{eq})}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\mathcal{I}_{4}(v,\omega,k).$ (85)
@@
$\hskip 28.45274pt+2\Gamma^{0}_{03}(\omega,\vec{k})\frac{1}{ik\tau_{eq}}\mathcal{I}_{1}(v,\omega,k)+\Gamma^{0}_{33}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,k)+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\frac{\mathcal{I}_{4}(v,\omega,k)}{i\tau_{eq}\omega}\bigg{\}},$
$\mathcal{E}_{3}\equiv-\tau_{eq}\int\,\frac{dp}{2\pi^{2}T_{0}}p^{2}\,p_{0}^{2}\,f_{eq}(p)\bigg{\{}\Gamma^{0}_{00}(\omega,\vec{k})\frac{1}{ik\tau_{eq}}\mathcal{I}_{1}(v,\omega,k)+\Gamma^{0}_{33}(\omega,\vec{k})\mathcal{I}_{5}(v,\omega,k)$
$\hskip 28.45274pt+2\Gamma^{0}_{03}(\omega,\vec{k})\frac{1-i\tau_{eq}\omega}{k^{2}\tau_{eq}^{2}}\mathcal{I}_{1}(v,\omega,k)+(\Gamma^{0}_{11}(\omega,\vec{k})+\Gamma^{0}_{22}(\omega,\vec{k}))\mathcal{I}_{6}(v,\omega,k)\bigg{\}},$ (86)
where $\mathcal{I}_{i}(v,\omega,k)$ functions for $(i=1,\cdots,9)$ are defined in below
$\mathcal{I}_{1}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\vec{v}(\vec{p})\cdot\vec{k}}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=1+\frac{i\,\omega\,\tau_{eq}-1}{2i\,\tau_{eq}\,k\,v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)},$
$\mathcal{I}_{2}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=-\frac{i}{k\,\tau_{eq}}\mathcal{I}_{1}(v,\omega,k),$
$\mathcal{I}_{3}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\,\omega\,v_{z}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\frac{i\,\omega(1-i\tau_{eq}\,\omega)}{k^{2}\,\tau_{eq}}\,\mathcal{I}_{1}(v,\omega,k),$
$\mathcal{I}_{4}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{i\,\tau_{eq}\,\omega\,v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=-\frac{i\tau_{eq}\omega}{2}\left(\frac{1-i\tau_{eq}\omega}{(\tau_{eq}\,k)^{2}}+\frac{(1-i\tau_{eq}\omega)^{2}+\tau_{eq}^{2}k^{2}v^{2}}{2v(i\tau_{eq}\,k)^{3}}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\right).$
$\mathcal{I}_{5}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{3}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\frac{(1-i\tau_{eq}\omega)^{2}}{(ik\tau_{eq})^{3}}+\frac{v^{2}}{3ik\tau_{eq}}-\frac{(1-i\tau_{eq}\omega)^{3}}{2vk^{4}\tau_{eq}^{4}}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)},$
$\mathcal{I}_{6}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\int\frac{d\Omega}{4\pi}\frac{v_{z}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{7}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{4}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{8}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{z}^{2}(\vec{p})v_{x}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}=\int\frac{d\Omega}{4\pi}\frac{v_{z}^{2}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$
@@
$\mathcal{I}_{9}(v,\omega,k)\equiv\int\frac{d\Omega}{4\pi}\frac{v_{x}^{2}(\vec{p})v_{y}^{2}(\vec{p})}{1+\tau_{eq}(-i\omega+i\vec{k}\cdot\vec{v}(\vec{p}))}$